OpenStudy (anonymous):

Let F(x) be the function 4x^2-7x+4. Then the quotient [F(6+H)-F(6)]/h can be simplified to for: Ah + B. How do I simplify this? My text book doesn't cover it... oddly enough Let F(x) be the function 4x^2-7x+4. Then the quotient [F(6+H)-F(6)]/h can be simplified to for: Ah + B. How do I simplify this? My text book doesn't cover it... oddly enough @Mathematics

OpenStudy (amistre64):

what are your outcomes for F(6) and F(6+h)

OpenStudy (amistre64):

and theres something missing between to and for

OpenStudy (anonymous):

Ah, Ah + B is suppose to be between to and for , then following For is What is A= and B =

OpenStudy (anonymous):

So i should just input 6+h and 6 into the polynomial? But how do I know what is Ah and what is B?

OpenStudy (amistre64):

A = 8 and B = -7 but thats what we should get to

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

I think its spose to read Ax + B

OpenStudy (anonymous):

different notation, I have seen it as Ax as well.

OpenStudy (amistre64):

4(6+h)^2-7(6+h)+4 - (4(6)^2-7(6)+4) 4(6+h)^2-7(6)+7(h)+4 - 4(6)^2+7(6)-4 4(6+h)^2+7(h) - 4(6)^2 4(36 + 12h +h^2)+7(h) - 4(6)^2 4(36) + 48h +4h^2+7h - 4(36) 48h +4h^2+7h ------------- = 48 + 7 + 4h h i spose it does work out to hs

OpenStudy (amistre64):

the 6 is what thru me off i think

OpenStudy (amistre64):

my math is off i think ... gotta dbl chk it

OpenStudy (anonymous):

After I do the computation how do i know which is Ax or B? Without knowing the rule, its ambiguous to me haha

OpenStudy (amistre64):

should be a -7

OpenStudy (amistre64):

the one with the h by it is the Ah part, and the one that stands alone is the B part

OpenStudy (amistre64):

just match the given pattern with what your results look like and equate the numbers

OpenStudy (amistre64):

yeah, i missed a sign up top; -*+ = - and i got +

OpenStudy (anonymous):

so 56 + 4h from the example you gave earlier a= 56 and B = 4? or am I jumping the gun

OpenStudy (amistre64):

48-7+4h is what it should be, i missed one of my signs in the process

OpenStudy (amistre64):

so we get 41 + 4h = Ah + B Ah = $h 41 = B

OpenStudy (amistre64):

cant type ....

OpenStudy (amistre64):

4h = Ah 41 = B so, A = 4 and B = 41

OpenStudy (anonymous):

ah okay cool, thanks man!